What is the molality of a 10% by mass NaCl solution when used in qualitative determination of tannins in plants?


1
Expert's answer
2021-10-03T18:15:02-0400

10% means 100 g of solution contains 10 g of NaCl


M = 58.44


n = 10/58.44 = 0.1711 mol


Density of the solution is 1.09


V = 100 / 1.09 = 91.74 mL = 0.09174 L


Molarity = 0.1711 / 0.09174 = 1.86 M